# thanks: Corresponding author

### v1

@@ Abstract
The rapid development of quantum computing technology has made it possible to study the thermodynamic properties of fermionic systems at finite temperatures through quantum simulations on a quantum computer. This provides a novel approach to the study of the chiral phase transition of fermionic systems. Among these, the quantum minimally entangled typical thermal states (QMETTS) algorithm has recently attracted considerable interest. The massive Thirring model, which exhibits a variety of phenomena at low temperatures, includes both a chiral phase transition and a topologically non-trivial ground state. It therefore raises the intriguing question of whether its phase transition can be studied using a quantum simulation approach. In this study, the chiral phase transition of the massive Thirring model and its dual topological phase transition are studied using the QMETTS algorithm. The results show that QMETTS is able to accurately reproduce the phase transition and thermodynamic properties of the massive Thirring model.
I Introduction
In recent years, there have been significant advances in the field of quantum computing technology [1]. It can be expected that the development of quantum computing will have a profound impact on the way to study physics in the future. Although the quantum computing is still in the noisy intermediate-scale quantum (NISQ) era [2, 3], the study of high-energy physics based on quantum computing algorithms [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], especially quantum simulations [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], is experiencing a period of rapid growth.
@@ I Introduction
Quantum simulation is also proposed to address the the case of many-body systems at finite temperatures, including approaches such as the quantum imaginary time evolution (QITE) algorithms. One of the QITE algorithms is the quantum minimally entangled typical thermal states (QMETTS) algorithm [34], which has recently been applied to the study of the chiral phase transition [35]. One of the advantages of quantum simulations is that, in contrast to numerical integration, they employ a physical approach, which is thought to enable them to bypass the sign problem encountered in numerical integration in traditional Monte Carlo methods, such as the notorious sign problem in lattice quantum chromodynamics (QCD) at finite chemical potentials [36, 37, 38]. The finite chemical potential is important in the study of phase transitions in QCD, because it plays a pivotal role in determining the location of the critical endpoint of the phase transition, which is of paramount importance in both experimental and theoretical contexts [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. In addition, there are other sign problems in the study of quark matter that arise under extreme conditions [53, 54, 55]. Therefore, the potential of quantum simulations to effectively study systems at finite temperatures is an avenue that warrants further investigation.
@@
In this work, we use the QMETTS algorithm to study the massive Thirring model [56]. The Thirring model is a toy model describing a self interacting fermion field, which has been extensively studied in previous works [57], including the study of the mass gap where quantum computing has already played a role [58]. One of the reasons for the importance of the Thirring model is that, it is exactly soluble in the massless case [59], while the massive case is soluble by the Bethe ansatz [60]. With the possibility of the exhibition of confinement, the Thirring model provides a chance to use a soluble interacting model to study the confinement [61, 62], which is a long-lasting subject in QCD without a final concluded description. The simultaneously broken of the chiral symmetry of the Thirring model is also an attractive phenomenon, which also plays an important role in the phase transition of QCD. Meanwhile, the Thirring model can be dual to the sine-Gordon theory [63] which contains topological structures, therefore the study of the Thirring model provides a simulation of topological soliton dynamics [64]. The massive Thirring model can also be dual to the Gross-Neveu (GN) model [65] with finite chemical potential, thus this work can also serve as another example of using quantum simulation to deal with systems with a finite chemical potential.
@@
Given the extensive prior research on the Thirring model, this study aims to ascertain whether the findings can be replicated using the QMETTS. This not only serves to validate the capabilities of the QMETTS, but also provides a detailed account of its practical applications. On the one hand, the QMETTS is better suited to the study of systems at high temperatures, whereas the massive Thirring model is also interesting in the low-temperature regime. On the other hand, the vacuum of the Thirring model is topologically non-trivial. It is well established that topologically non-trivial ground states cannot be connected to trivial ones by quantum circuits of finite depth in the thermodynamic limit [66]. However, in finite-volume systems, the question of whether the barriers between different topological sectors can be crossed by QMETTS is another intriguing topic for exploration.
@@
The remainder of the paper is organized as follows. In Section II, a brief introduction to the massive thirring model is presented. The QMETTS algorithm is discussed in Section III. Section IV presents numerical results on the chiral condensation and topological charge obtained by using the QMETTS algorithm. Section V is a summary of the conclusions.
II A brief introduction of the massive Thirring model
@@
where $\sigma^{i}(n)$ are Pauli matrices. In order to study the chiral condensation, we use the staggered fermion which keeps the chiral symmetry in the massless limit [67, 68, 69]. For a staggered fermion, the components of a spinor are distributed within two sites. Denoting $h=2n$ as the even sites, the Dirac fermion field $\psi(h)$ can be written as,
@@
In this work, the Minkowski and Euclidean cases of interaction are both considered. The Hamiltonian in these cases are denoted as $H_{M}$ and $H_{E}$, respectively, which can be written as,
$\begin{split}&H_{E,M}=a\sum_{h}\left\{\bar{\psi}(h)\left(i\gamma^{1}\partial_{1}+m\right)\psi(h)\right.\\ &\left.-\frac{g^{2}}{4}\left[\left(\bar{\psi}(h)\gamma^{0}\psi(h)\right)^{2}\pm\left(\bar{\psi}(h)\gamma^{1}\psi(h)\right)^{2}\right]\right\}.\end{split}$ (4)
For the derivation of $\psi$, we use forward derivation and backward derivation for different components of the spinor,
$\begin{split}&\partial_{h}\psi(h)=\frac{1}{a\sqrt{a}}\begin{pmatrix}\chi(h+2)-\chi(h)\\ \chi(h+1)-\chi(h-1)\end{pmatrix}.\end{split}$ (5)
The staggered fermion is then transformed to the Jordan-Wigner representation [70],
@@
where $\sigma^{i}(n)$ are Pauli matrices sitting on sites, note that different from the $\gamma$ matrices in Eq. (2), Pauli matrices in Eq. (6) are not in the spinor space.
When the total number of sites $N$ is even, one is able to verify the following equivalences,
$\begin{split}&a\sum_{h}\bar{\psi}(h){\rm i}\gamma_{1}\partial_{x}\psi(h)=\\ &\frac{1}{2a}\sum_{i=x,y}\left\{\sum_{n}^{N-2}\sigma^{i}(n)\sigma^{i}(n+1)\right.\\ &\left.+(-1)^{\frac{N}{2}}\sigma^{i}(0)\sigma^{i}(N-1)\prod_{j=1}^{N-2}\sigma^{z}(j)\right\},\end{split}$ (7)
$\begin{split}&a\sum_{h}\bar{\psi}(h)\psi(h)=\sum_{n=0}(-1)^{n}\frac{1+\sigma^{z}(n)}{2},\end{split}$ (8)
$\begin{split}&a\sum_{h}\left(\bar{\psi}(h)\gamma_{0}\psi(h)\right)^{2}\\ &=\frac{1}{a}\sum_{n}\left(\frac{3}{4}+\sigma^{z}(n)+\frac{\sigma^{z}(n)\sigma^{z}(n+1)}{4}\right),\end{split}$ (9)
$\begin{split}&a\sum_{h}\left(\bar{\psi}(h)\gamma_{1}\psi(h)\right)^{2}=-\frac{1}{4a}\sum_{n}\left(1-\sigma^{z}(n)\sigma^{z}(n+1)\right),\end{split}$ (10)
$\begin{split}&a\sum_{h}\bar{\psi}(h)\gamma_{0}\psi(h)=\sum_{n=0}\frac{1+\sigma^{z}(n)}{2},\end{split}$ (11)
where the terms with higher orders of $a$ are ignored to keep the translational invariance of the Hamiltonian. The ignored terms are shown in Appendix A. Using,
$\begin{split}&a\sum_{h}(\bar{\psi}(h)\psi(h))^{2}=\frac{1}{4a}\sum_{n}\left(1+\sigma^{z}(n)\sigma^{z}(n+1)\right),\\ \end{split}$ (12)
the massive Thirring model in $d=1+1$ can be related to the GN model with chemical potential, whose Hamiltonian is,
$\begin{split}&H_{GN}(m,g,\mu)=\bar{\psi}\left(i\gamma_{1}\partial_{1}+m\right)\psi\\ &-g(\bar{\psi}\psi)^{2}-\mu\bar{\psi}\gamma_{0}\psi,\end{split}$ (13)
where $\mu$ is chemical potential. Ignoring the constants and $\mathcal{O}(a)$ terms, it can be seen that,
$\begin{split}&H_{M}(m,g)=H_{GN}(m,0,\frac{ag^{2}}{2})\\ &H_{E}(m,g)=H_{GN}(m,\frac{g^{2}}{2},\frac{ag^{2}}{2})\\ \end{split}$ (14)
It has been shown in Ref. [35] by using QMETTS that, there is a chiral phase transition where the chiral symmetry is spontaneously broken. Therefore, it can be expected that the massive Thirring model will also present a chiral phase transition. Furthermore, since the interaction plays a role as the chemical potential, the spontaneous symmetry breaking is induced by the interaction in the massive Thirring model without a chemical potential. To investigate such a phenomenon, the chiral condensation as a function of $g$ and temperature $T$ is considered.
@@
III A brief introduction of the QMETTS algorithm
QMETTS is a promising tool for studying the thermal properties of quantum systems which leverages the concept of quantum typicality. QMETTS computes the typical thermal states, which are pure states that serve as an efficient approximation for simulating the thermal properties of quantum systems. The idea of QMETTS is to approximate $e^{-\beta H}|\phi\rangle$ using $e^{-iA\beta}|\phi\rangle$, where $A$ is an operator which will be introduced later. Then the trace ${\rm tr}\left[e^{-\beta H}\right]$ can be calculated using a Hilbert space $|\phi_{i}\rangle$ as ${\rm tr}\left[e^{-\beta H}\right]=\sum_{i}\langle\phi_{i}|e^{-\beta H}|\phi_{i}\rangle$. Note that the trace can also be calculated in a stochastic approach using a set of randomly generated states if the Hilbert space is too large [72].
We briefly introduce the QMETTS following Ref. [35]. For an arbitrary state $\phi$, assume,
$\begin{split}&e^{-iA\Delta\beta}|\phi\rangle\approx\sqrt{C(\Delta\beta)}e^{-\Delta\beta H}|\phi\rangle,\\ \end{split}$ (15)
@@ III A brief introduction of the QMETTS algorithm
$\begin{split}&C(\Delta\beta)=\langle\phi|e^{-2\Delta\beta H}|\phi\rangle\approx 1-2\Delta\beta\langle\phi|H|\phi\rangle,\\ \end{split}$ (16)
@@
Denoting $j=l_{0}+l_{1}\times 4+l_{2}\times 4^{2}+\ldots$ as a combined index of $\left(l_{0},l_{1},l_{2},\ldots\right)$ with $l_{i}$ the integers satisfying $0\leq l_{i}\leq 3$, an Hermitian operator $A$ on $N$ sites (or $N$ qubits) can be expanded as,
$\begin{split}&A=\sum_{j}a_{j}\hat{\sigma}_{j}=\sum_{j}a_{j}\prod_{n}^{N}\sigma^{l_{n}}(n),\\ \end{split}$ (17)
with $a_{j}$ are real coefficients, and $\sigma^{0,1,2,3}=1$, $\sigma^{x}$, $\sigma^{t}$, and $\sigma^{z}$, respectively. It is found that, Eq. (15) is a well approximation when $\Delta\beta$ is small, and $\{a_{j}\}$ is the solution of the equation,
$\begin{split}&\left(S+S^{T}\right)a=b,\end{split}$ (18)
@@
$\begin{split}&S_{j_{1},j_{2}}=\langle\phi|\hat{\sigma}_{j_{i}}\hat{\sigma}_{j_{2}}|\phi\rangle,\\ &b_{j}=\frac{-{\rm i}}{\sqrt{C(\Delta\beta)}}\langle\phi|\left(H\hat{\sigma}_{j}-\hat{\sigma}_{j}H\right)|\phi\rangle,\\ \end{split}$ (19)
where both $S$ and $b$ can be obtained by measurements. The details of the measurements are shown in Appendix B. The matrix elements of $S$ can also be obtained by using the Hadamard test.
As will be introduced later, when considering an observable at $\beta$, it only needs to calculate $\phi(\beta/2)$. In this work, we divide $\beta/2$ by $K$ steps, i.e., $\Delta\beta=\beta/2K$. Denoting $C(\Delta\beta,\beta^{\prime})=1-2\Delta\beta\langle\phi(\beta^{\prime})|H|\phi(\beta^{\prime})\rangle$, and $A(a(\Delta\beta,\beta^{\prime}))$ as the operator $A$ with coefficients $a_{j}$ solved by using $\phi(\beta^{\prime})$, $\phi(\beta/2)$ can be obtained iteratively as,
$\begin{split}&\left|\phi\left(\frac{\beta}{2}\right)\right\rangle=\prod_{k=0}^{K-1}e^{-{\rm i}A\left(a(\Delta\beta,k\Delta\beta)\right)\Delta\beta}|\phi(0)\rangle,\\ &C_{\phi}=\prod_{k=0}^{K-1}C(\Delta\beta,k\Delta\beta),\\ &\left|\Phi\left(\frac{\beta}{2}\right)\right\rangle=\frac{1}{\sqrt{C_{\phi}}}\left|\phi\left(\frac{\beta}{2}\right)\right\rangle.\end{split}$ (20)
The $e^{-{\rm i}A\left(a(\Delta\beta,k\Delta\beta)\right)\Delta\beta}|\phi(k\Delta\beta)\rangle$ can be evaluated using Trotter decomposition with $t$ steps [73].
Then, for an observable $\hat{O}$ in interest,
@@
$\begin{split}&{\rm tr}\left(e^{-\beta H}\hat{O}\right)=\sum_{i}\left\langle\Phi_{i}\left(\frac{\beta}{2}\right)\right|\hat{O}\left|\Phi_{i}\left(\frac{\beta}{2}\right)\right\rangle,\\ &{\rm tr}\left(e^{-\beta H}\right)=\sum_{i}\left\langle\Phi_{i}\left(\frac{\beta}{2}\right)|\Phi_{i}\left(\frac{\beta}{2}\right)\right\rangle.\end{split}$ (22)
@@
IV Numerical results
In this work, $N=4$ is used. According to Eqs. (4) and (7-10), the Hamiltonian can be written as,
$\begin{split}&aH_{E}=h_{1}+amh_{2}-\frac{1}{4}g^{2}h_{3}-\frac{1}{4}g^{2}h_{4},\\ &aH_{M}=h_{1}+amh_{2}-\frac{1}{4}g^{2}h_{3}+\frac{1}{4}g^{2}h_{4},\\ \end{split}$ (25)

### v2

# Quantum simulation of the phase transition of the massive Thirring model

@@ Abstract
Recent advancements in quantum computing technology have enabled the study of fermionic systems at finite temperature via quantum simulations. This presents a novel approach to investigating the chiral phase transition in such systems. Among these, the quantum minimally entangled typical thermal states (QMETTS) algorithm has recently attracted considerable interest. The massive Thirring model, which exhibits a variety of phenomena at low temperatures, includes both a chiral phase transition and a topologically non-trivial ground state. It therefore raises the intriguing question of whether its phase transition can be studied using a quantum simulation approach. In this study, the chiral phase transition of the massive Thirring model and its dual topological phase transition are studied using the QMETTS algorithm. Numerical results are obtained on a classical computer simulating circuit-based quantum computations. The results show that QMETTS is able to accurately reproduce the phase transition and thermodynamic properties of the massive Thirring model.
1 Introduction
In recent years, significant advances have been made in quantum computing technology Arute:2019zxq . It is expected that the development of quantum computing will profoundly impact the study of physics in the future. Although quantum computing remains in the noisy intermediate-scale quantum (NISQ) era Preskill:2018jim ; Chen:2021num , research in high energy physics using quantum algorithms Fang:2024ple ; Scott:2024txs ; Roggero:2018hrn ; Roggero:2019myu ; Atas:2021ext ; Lamm:2019uyc ; Li:2021kcs ; Perez-Salinas:2020nem ; Jordan:2011ci ; Mueller:2019qqj ; Guan:2020bdl ; Wu:2021xsj ; Zhang:2023ykh ; Zhang:2024ebl ; Zhu:2024own ; Fadol:2022umw ; Wu:2020cye ; Terashi:2020wfi ; Yang:2024bqw , particularly quantum simulations Feynman:1981tf ; Georgescu:2013oza ; Bauer:2022hpo ; Carena:2022kpg ; Gustafson:2022xdt ; Lamm:2024jnl ; Carena:2024dzu ; Li:2023vwx ; Cui:2019sfz ; Zou:2021pvl ; Echevarria:2020wct , is rapidly advancing.
@@ 1 Introduction
Quantum simulation has also been proposed to address many-body systems at finite temperature, using methods such as quantum imaginary time evolution (QITE) algorithms. One notable QITE-based method is the Quantum Minimally Entangled Typical Thermal States (QMETTS) algorithm Motta:2019yya , which has recently been applied to the study of the chiral phase transition Czajka:2021yll . One advantage of quantum simulations is that unlike numerical integration, they use a physical approach. As a result, quantum simulations are believed to circumvent the sign problems found in traditional Monte Carlo methods, including the notorious case in lattice quantum chromodynamics (QCD) at finite chemical potential Hasenfratz:1983ba ; deForcrand:2009zkb ; Gattringer:2016kco . The finite chemical potential is important in the study of phase transitions in QCD, because it plays a crucial role in determining the location of the critical endpoint of the phase transition, which is of paramount importance in both experimental and theoretical contexts Fukushima:2013rx ; Busza:2018rrf ; STAR:2010vob ; Gupta:2011wh ; Andronic:2017pug ; Luo:2020pef ; STAR:2021iop ; Stephans:2006tg ; NA49-future:2006qne ; Gazdzicki:2008kk ; Luo:2017faz ; Stephanov:2004wx ; Fodor:2004nz ; Gavai:2004sd . In addition, there are other sign problems in the study of quark matter that arise under extreme conditions Yang:2022zob ; Yang:2023vsw ; Yang:2023zzx . Therefore, the potential of quantum simulations to effectively study systems at finite temperatures is an avenue that warrants further investigation.
@@
In this work, we use the QMETTS algorithm to study the massive Thirring model Thirring:1958in . The Thirring model is a toy model describing a self interacting fermion field, which has been extensively studied in previous works Sachs:1995dm , including the study of the mass gap where quantum computing has already played a role Mishra:2019xbh . One of the reasons for the importance of the Thirring model is that it is exactly soluble in the massless case Hagen1967 , while the massive case is soluble by the Bethe ansatz Korepin_Bogoliubov_Izergin_1993 . With the possibility of the exhibition of confinement, the Thirring model provides a chance to use a soluble interacting model to study the confinement Alvarez-Estrada:1997vmv ; PintoBarros:2018bzz , which still lacks a complete theoretical description. The simultaneous breaking of the chiral symmetry in the Thirring model is another attractive phenomenon, as it mirrors key features of QCD phase transitions. Furthermore, the Thirring model is dual to the sine-Gordon theory Coleman:1974bu which possesses topological structures. This duality enables the simulation of topological soliton dynamics through the Thirring model Banuls:2017evv . As will be shown, the massive Thirring model can also be dual to the Gross-Neveu (GN) model Gross:1974jv with a finite chemical potential, thus this work can also serve as another example of using quantum simulation to study a system with a finite chemical potential.
@@
Given the extensive prior research, our aim is to test whether QMETTS can accurately reproduce known results of the Thirring model. This not only serves to validate the capabilities of the QMETTS, but also provides a detailed account of its practical applications. On the one hand, the QMETTS is better suited to the study of systems at high temperatures, whereas the massive Thirring model is also interesting in the low-temperature regime. On the other hand, the vacuum of the Thirring model is topologically non-trivial. It is well established that topologically non-trivial ground states cannot be connected to trivial ones by quantum circuits of finite depth in the thermodynamic limit Chen:2010gda . However, in finite-volume systems, the question of whether the barriers between different topological sectors can be crossed by QMETTS is another intriguing topic for exploration.
@@
The remainder of this paper is organized as follows. In Section 2, a brief introduction to the massive thirring model is presented. The QMETTS algorithm is discussed in Section 3. Section 4 presents numerical results on the chiral condensation and topological charge obtained by using the QMETTS algorithm. In section 5, we present discussions on the errors. Section 6 is a summary of the conclusions.
2 A brief introduction of the massive Thirring model
@@
where $\sigma^{i}$ are Pauli matrices. In order to study the chiral condensation, we use the staggered fermion which preserves the chiral symmetry in the massless limit Kogut:1974ag ; Kluberg-Stern:1983lmr ; Morel:1984di . For a staggered fermion, the components of a spinor are distributed within two sites. Let $h=2n$ denote the even sites, the Dirac fermion field $\psi(h)$ is then written as,
@@
In this work, the Minkowski and Euclidean cases of interactions are both considered. The Hamiltonian in these cases are denoted by $H_{M}$ and $H_{E}$, respectively, and are given by,
$\begin{split}&H_{E,M}=a\sum_{h}\left\{\bar{\psi}(h)\left(i\gamma^{1}\partial_{1}+m\right)\psi(h)-\frac{g^{2}}{4}\left[\left(\bar{\psi}(h)\gamma^{0}\psi(h)\right)^{2}\pm\left(\bar{\psi}(h)\gamma^{1}\psi(h)\right)^{2}\right]\right\}.\end{split}$ (4)
To discretize the derivative of $\psi$, forward and backward finite differences are used for different spinor components,
[…]
The staggered fermion is then transformed into the Jordan-Wigner representation JW ,
@@
where $\sigma^{i}(n)$ are Pauli matrices sitting on sites, note that different from the $\gamma$ matrices in Eq. (2), Pauli matrices in Eq. (6) do not act in spinor space.
When the total number of sites $N$ is even, the following equivalences can be verified,
@@
where the terms with higher orders of $a$ are neglected to keep the translational invariance of the Hamiltonian. The omitted terms are shown in Appendix A. Using,
[…]
the massive Thirring model in $d=1+1$ can be related to the GN model with a chemical potential, whose Hamiltonian is,
$\begin{split}&H_{GN}(m,g,\mu)=\bar{\psi}\left(i\gamma_{1}\partial_{1}+m\right)\psi-g(\bar{\psi}\psi)^{2}-\mu\bar{\psi}\gamma_{0}\psi,\end{split}$ (9)
where $\mu$ is chemical potential. Neglecting the constants and $\mathcal{O}(a)$ terms, it can be seen that,
[…]
It has been shown in Ref. Czajka:2021yll by using QMETTS that there is a chiral phase transition where the chiral symmetry is spontaneously broken. Therefore, it can be expected that the massive Thirring model will also present a chiral phase transition. Furthermore, since the interaction effectively acts as a chemical potential, it induces spontaneous symmetry breaking in the massive Thirring model even without an explicit chemical potential. To investigate such a phenomenon, the chiral condensation as a function of $g$ and temperature $T$ is considered.
The duality between the sine-Gordon and the Thirring model has been proved for zero temperature Coleman:1974bu and finite temperature Delepine:1997bz . It has been established that $\bar{\psi}\gamma^{\mu}\psi$ corresponds to $\left(\beta/2\pi\right)\epsilon^{\mu\nu}\partial_{\nu}\phi$, where $\phi$ is the scalar field in the sine-Gordon model. Note that $\left(\beta/2\pi\right)\int dx\partial_{x}\phi$ is a topological charge measuring the winding number in $\phi$. Consequently, the fermion number $\langle\bar{\psi}\gamma^{0}\psi\rangle$ corresponds to a topological charge in the sine-Gordon model. Therefore, $\langle\bar{\psi}\gamma^{0}\psi\rangle$ serves as another key observable in this study.
3 A brief introduction of the QMETTS algorithm
QMETTS is a promising tool for studying the thermal properties of quantum systems, leveraging the concept of quantum typicality. QMETTS computes the typical thermal states, which are pure states that serve as an efficient approximation for simulating the thermal properties of quantum systems. The idea of QMETTS is to approximate $e^{-\beta H}|\phi\rangle$ using $e^{-ia^{-1}\beta A}|\phi\rangle$, where $A$ is an operator which will be introduced later. Then the trace ${\rm tr}\left[e^{-\beta H}\right]$ can be calculated using a Hilbert space $|\phi_{i}\rangle$ as ${\rm tr}\left[e^{-\beta H}\right]=\sum_{i}\langle\phi_{i}|e^{-\beta H}|\phi_{i}\rangle$. Note that the trace can also be calculated in a stochastic approach using a set of randomly generated states if the Hilbert space is too large Shen:2024fcj .
We briefly introduce the QMETTS following Ref. Czajka:2021yll . For an arbitrary state $\phi$, assume,
[…]
@@ 3 A brief introduction of the QMETTS algorithm
[…]
@@
Denoting $j=l_{0}+l_{1}\times 4+l_{2}\times 4^{2}+\ldots$ as a combined index of $\left(l_{0},l_{1},l_{2},\ldots\right)$ with $l_{i}$ the integers satisfying $0\leq l_{i}\leq 3$, any Hermitian operator (in our case the $A$ operator) on $N$ sites (or $N$ qubits) can be expanded as,
[…]
where the $c_{j}$ are real coefficients, and $\sigma^{0,1,2,3}=1$, $\sigma^{x}$, $\sigma^{t}$, and $\sigma^{z}$, respectively. It is found that, Eq. (11) provides a good approximation when $\Delta\beta$ is small, and $\{c_{j}\}$ is the solution of the equation,
$\begin{split}&\left(S+S^{T}\right)c=b,\end{split}$ (14)
@@
[…]
where both $S$ and $b$ can be obtained by measurements or Hadamard tests. The details of the measurements are shown in Appendix B.
As will be introduced later, when considering an observable at $\beta$, it only needs to calculate $\phi(\beta/2)$. In this work, we divide $\beta/2$ by $K$ steps, i.e., $\Delta\beta=\beta/2K$. Denoting $C(\Delta\beta,\beta^{\prime})=1-2\Delta\beta\langle\phi(\beta^{\prime})|H|\phi(\beta^{\prime})\rangle$, and $A(c_{j}(\Delta\beta,\beta^{\prime}))$ as the operator $A$ with coefficients $c_{j}$ solved by using $\phi(\beta^{\prime})$, $\phi(\beta/2)$ can be obtained iteratively as,
$\begin{split}\left|\phi\left(\frac{\beta}{2}\right)\right\rangle&=\prod_{k=0}^{K-1}e^{-{\rm i}A\left(c_{j}(\Delta\beta,k\Delta\beta)\right)a^{-1}\Delta\beta}|\phi(0)\rangle,\\ C_{\phi}&=\prod_{k=0}^{K-1}C(\Delta\beta,k\Delta\beta),\\ \left|\Phi\left(\frac{\beta}{2}\right)\right\rangle&=\frac{1}{\sqrt{C_{\phi}}}\left|\phi\left(\frac{\beta}{2}\right)\right\rangle.\end{split}$ (16)
Each $e^{-{\rm i}A\left(c_{j}(\Delta\beta,k\Delta\beta)\right)a^{-1}\Delta\beta}|\phi(k\Delta\beta)\rangle$ can be evaluated using a Trotter decomposition with $t$ steps Lloyd:1996aai .
Then, for an observable $\hat{O}$ of interest,
@@
$\begin{split}{\rm tr}\left(e^{-\beta H}\hat{O}\right)&=\sum_{i}\left\langle\Phi_{i}\left(\frac{\beta}{2}\right)\right|\hat{O}\left|\Phi_{i}\left(\frac{\beta}{2}\right)\right\rangle,\\ {\rm tr}\left(e^{-\beta H}\right)&=\sum_{i}\left\langle\Phi_{i}\left(\frac{\beta}{2}\right)|\Phi_{i}\left(\frac{\beta}{2}\right)\right\rangle.\end{split}$ (18)
@@
4 Numerical results
In this work, $N=4$ is used. According to Eqs. (4) and (7), the Hamiltonian can be expressed as,
[…]